# Resonant Solitary States in Complex Networks

### v1

@@ Abstract
We now apply our result to the complex network shown in Figs. 2 and S2. Solitary states with intermediate frequency are located at leaves with neighbor degree $\geq$ 6, (cf. Nitzbon et al. (2017)). Figure 4 shows the solutions of Eq. (11) for $z=74$ for slightly heterogeneous parameters.
@@
As many network modes are excited, the response curve is considerably more complex. However, simulation results reveal that there is still one dominant mode with high localization at the root node $k$, which is in close resonance with the solitary frequency, and which occurs frequently from random initial conditions.
There are several tentative observations we can make by looking at these examples, (Figs. S4 and S5)): (i) Eq. (11) tends to underestimate the magnitude of the most prominent solution for $\omega_{s}$ slightly, this can already be observed in the effective model; (ii) pairs of close stable and unstable solutions, which are caused by narrow peaks, result in solitary states with a small basin of attraction; (iii) pairs close to their annihilating bifurcation, caused by peaks that barely exceed the straight line at the intersection, also cause small basins; (iv) the most prominent solitary state lies at a frequency that has a high and broad peak and some distance to the next solutions.
Leaf nodes with high neighbor degree Nitzbon et al. (2017), typically produce category (iv) situations (SM I.A). Thus, these findings explain why such nodes feature solitary states with intermediate frequencies.
@@
The derivation of the minimal model, and the explicit relationship between network mode and neighbor degree (SM I.A and II.A), suggests that the neighbor degree can serve as an upper bound of the intermediate frequency in the complex network case. This is also observed numerically (Fig. S8).
@@
Conclusion.— We present an analytical description of the effective coupling between a solitary oscillator and the synchronized cluster in a complex network. To approximate the trajectories, we utilize partial linearization and an averaging theorem. This way, we discover that a resonant excitation of the linear modes of the synchronized cluster can couple coherently to the shifted frequency of the solitary state, resulting in a large energy flow between the synchronized cluster and the solitary oscillator and effectively shifting the frequency.
@@
Our framework predicts the location of solitary oscillations in reduced models, and is in excellent agreement with what we see in complex networks. With this, we can explain that solitary states appear mostly in specific parts of the network Berner et al. (2021), more specifically leaves and tree-shaped structures, which act as weak points in perturbation scenarios Menck et al. (2014); Nitzbon et al. (2017); Halekotte and Feudel (2020); Halekotte et al. (2021). These results have important implications for improving grid stability through modifications of parameters Motter et al. (2013); Wassmer et al. (2021); Menara et al. (2022), topology Witthaut and Timme (2012); Menck et al. (2014); Schultz et al. (2016); Kaiser et al. (2021), and targeted control Taher et al. (2019).
@@
While the presented mechanism can explain the potential existence and stability of intermediate solitary states, there are still open questions that require further investigation. A key practical question is the estimation of the basin boundaries of these states, which would provide deeper insights into the dynamics and robustness of synchronous systems, against perturbations Manik et al. (2014); Klinshov et al. (2015); Halekotte and Feudel (2020); Halekotte et al. (2021); Munyayev et al. (2022).
@@
Detailed numerical studies suggest that additional classes of attractive states exist, and the interaction of losses on the lines and resonant energy flow remains unexplored Hellmann et al. (2020). Furthermore, a more detailed study of the interaction of multiple solitary nodes with distinct frequencies or in frequency clusters is needed.
Finally, while the method introduced above focuses on phase oscillators, it would be interesting to extend it to phase-amplitude oscillators that play a central role in networks such as power grids Kogler et al. (2022). We expect that the overall analytical approach we introduced can be further developed and adapted to address these types of questions.
@@
The code for reproducing our simulations and figures can be found in Niehues (2023). The repository can also be applied to find solitary frequencies in complex networks by solving Eq. (11).
@@ Acknowledgements.
Acknowledgments.— Initial investigations of this phenomenon grew out of work by Anton Plietzsch and discussions with Paul Schultz, Chris Bick, Jeroen Lamb, Deniz Eroglu, Tiago Pereira, and Carsten Grabow during a stay at Nesim Mathematics Village. Part of this work was funded by the DFG grant CoCoHype (DFG KU 837/39-2, J.N., J.K., F.H., M.A.). J.N. gratefully acknowledges support by the Studienstiftung des Deutschen Volkes scholarship foundation and Berlin International Graduate School in Model and Simulation based Research (BIMoS) at Technische Universität Berlin.

### v3

@@ VI.2 Example II: Complex network
We now apply our result to the complex network shown in Figs. 2 and S2. For models of this complexity, the self-consistent equation has to be solved numerically for every node. We observe that intermediate resonant solitary states are located at leaf nodes with neighbor degree $\geq$ 6, (cf. Nitzbon et al. (2017)). Figure 4 shows the solutions of Eq. (21) for $z=74$ for slightly heterogeneous parameters.
@@
As many network modes are excited, the response curve $p_{s}(\omega_{s})$ is considerably more complex than for the two-node model. However, simulation results reveal that there is still one dominant mode with high localization at the root node $k$. Solitary states obtained from random initial conditions are frequently found in close resonance with the dominant mode, or close to the natural frequency.
There are several tentative observations we can make by looking at complex networks, (see also SM II.C, Figs. S5 and S6 for more examples): (i) If the self-consistent equation yields more solitary states than the one at the natural frequency, there are two dominating solitary states: the one at the natural frequency, and one in resonance with a localized network mode. There can also be many additional solutions with small basins of attraction, cf. Figs. S4 and S5.
(ii) Eq. (21) tends to overestimate the magnitude of the most prominent solution for $\omega_{s}$ slightly, which can already be observed in the effective model (cf. Figs. 3 and 4). (ii) the most prominent natural solitary state lies at an intermediate frequency that has a high and broad peak in $p_{s}$ and some distance to the next solutions. This can be understood by considering that every peak in $p_{s}$ produces a pair of solutions, a stable one and an unstable one, and that the solitary frequency $\varphi_{z}(t)$ is modulated around $\omega_{s}$, cf. Fig. 1. Leaf nodes with high neighbor degree Nitzbon et al. (2017), typically produce such situations (cf. SM I.A) with a highly localized network mode that results in a high peak. Thus, these findings explain why such nodes feature resonant solitary states with intermediate frequencies. (iii) Besides the main peak at the most localized network mode, there can be several minor peaks that result in solitary states with much smaller basins of attraction due to three reasons: First, the peaks might be narrow, such that the stable and unstable solution are close, second, the peak might not exceed the intersection with the straight line by much, hence the pair of solutions is already close to their annihilation bifurcation, and third, the peak might be close to other peaks and their solutions. This can account for the observation that parts of the phase space can be dominated by basin boundaries of many, different, unlikely states Gelbrecht et al. (2020), inducing transient chaos Halekotte et al. (2021).
@@
As a closing remark, the derivation of the minimal model, and the explicit relationship between network mode and neighbor degree (SM I.A and II.A), suggest the following. Even in complex networks, the neighbor degree $d_{k}$ can serve as an upper bound for the (squared) intermediate solitary frequency, when weighted with the coupling $\kappa_{ij}$. This is also observed numerically (Fig. S8).
VII Summary and discussion
@@
We present an analytical description of the effective coupling between a solitary oscillator and the synchronized cluster in a complex network. To approximate the trajectories, we utilize partial linearization and an averaging theorem. This way, we discover that a resonant excitation of the linear modes of the synchronized cluster can couple coherently to the shifted frequency of the solitary state, resulting in a large energy flow between the synchronized cluster and the solitary oscillator, and effectively shifting the frequency.
We have mainly found that intermediate solitary states, first observed in Nitzbon et al. (2017), are actually resonant solitary states, and we have uncovered how the network’s topology shapes their properties. Furthermore, we have unified their analytical description with the one of natural solitary states Menck et al. (2014), and shown that there is a spectrum connecting those extremes.
@@
Our self-consistent framework predicts solitary states in reduced models, where it is solvable, and is also in excellent agreement with what we observe numerically in the original complex networks. With this structure-function relationship, we can explain the observation that solitary states appear mostly in specific parts of the network Berner et al. (2021), more specifically leaf nodes and tree-shaped structures, which act as weak points in perturbation scenarios Menck et al. (2014); Nitzbon et al. (2017); Halekotte and Feudel (2020); Halekotte et al. (2021). These results have important implications for improving grid stability through modifications of parameters Motter et al. (2013); Wassmer et al. (2021); Menara et al. (2022) and topology Witthaut and Timme (2012); Menck et al. (2014); Schultz et al. (2016); Kaiser et al. (2021): Our framework can be used to evaluate how these modifications affect the presence and stability of potentially harmful solitary attractors and their proximity to the synchronous state. Furthermore, we can efficiently identify the troublemaker nodes that tend to desynchronize the easiest due to their intermediate frequencies: resonant solitary states at certain leaf nodes (dense sprout nodes, see Sup ; Nitzbon et al. (2017)). We can identify them from the network topology only, and without expensive simulations. It has been shown that targeted control of such troublemaker nodes can improve grid stability drastically Taher et al. (2019).
VIII Conclusion and Outlook
@@
While the presented mechanism can explain the potential existence and stability of resonant solitary states, there are still open questions that require further investigation. A key practical question is the estimation of the basin boundaries of these states, which would provide deeper insights into the dynamics and robustness of synchronous systems against perturbations Manik et al. (2014); Klinshov et al. (2015); Halekotte and Feudel (2020); Halekotte et al. (2021); Munyayev et al. (2022). It is known that the basin boundaries depend on the phase and frequency Menck et al. (2014); Hellmann et al. (2020); Halekotte et al. (2021). Averaging the phase dependence out, an estimate of the basin boundaries in $\omega$-direction could be estimated by the turning points of $Z(\omega)$ around a stable solution.
@@
Detailed numerical studies suggest that additional classes of attractive states exist, and the interaction of losses on the lines and resonant energy flow remains unexplored Hellmann et al. (2020). Furthermore, a more detailed study of the interaction of multiple solitary nodes with distinct frequencies or in frequency clusters is needed (see SM I.F for an outline).
Finally, while this paper focuses on explaining a known phenomenon of Kuramoto oscillators, which exhibit a purely sinusoidal coupling, our derivation generally only demands a $2\pi$-periodic coupling function. This includes coupling functions with phase shifts and harmonics. It would also be interesting to treat phase-amplitude oscillators that play a central role in networks such as power grids Kogler et al. (2022). We expect that the overall analytical approach we introduced is a promising tool in describing the synchronization of rotation and oscillation, and leave it to future work to develop and adapt our theory to address a wide range of models.
Code Availability
@@
The numerical experiments were conducted with the PowerDynamics.jl package Plietzsch et al. (2022). Further details can be found in SM III. The code for reproducing our simulations and figures can be found in Niehues (2023). The repository can also be applied to find solitary frequencies in complex networks by solving Eq. (21).
@@ Acknowledgements.
Acknowledgments
Initial investigations of this phenomenon grew out of work by Anton Plietzsch and discussions with Paul Schultz, Chris Bick, Jeroen Lamb, Deniz Eroglu, Tiago Pereira, and Carsten Grabow during a stay at Nesim Mathematics Village.
Part of this work was funded by the DFG grant CoCoHype (DFG KU 837/39-2; J.N., J.K., F.H., M.A.) and BMWK grant OpPoDyn (03EI1071A; J.N., F.H.).
J.N. gratefully acknowledges support by Studienstiftung des Deutschen Volkes scholarship foundation and Berlin International Graduate School in Model and Simulation based Research (BIMoS) at Technische Universität Berlin. S.Y. and J.K. acknowledge funding by DFG, German Research Foundation, Project No. 411803875.